OpenStudy (anonymous):

A meter stick is suspended vertically at a pivot point 28 cm from the top end. It is rotated on the pivot until it is horizontal and then released from rest. What will be its maximum angular velocity (in radians/second)? Please help. I almost feel like I can't solve this at all because of lack of information, but I know its hidden in the problem. Any help is appreciated.

OpenStudy (amistre64):



OpenStudy (amistre64):

just a stab at it ... correctness is not assumed i takes .72m = 9.8 t^2/2 2(.72)/9.8 = t^2 sqrt(2(.72)/9.8) = t i assume is the time it takes to reach the bottom of its arc (its fastest point)

OpenStudy (irishboy123):